OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!! What is the eccentricity of the ellipse given by the equation ? 1.59 1.24 0.81 0.63

OpenStudy (magbak):

\[(x^2/49 )+ (y^2/81)=1\]

OpenStudy (magbak):

@Hero @satellite73

OpenStudy (magbak):

@Mertsj

OpenStudy (magbak):

@Loser66

OpenStudy (mertsj):

a^2-b^2=c^2 Find c. eccentricity is c/a

OpenStudy (magbak):

Help me find the answer PLEASE/

OpenStudy (loser66):

first, find c^2 , follow Mertsj 's instruction, your a^2 =81, b^2 =49, c^2=?

OpenStudy (magbak):

OK can you wait till I find the answer.

OpenStudy (magbak):

@Mertsj

OpenStudy (magbak):

Is the entricity 0.395061728

OpenStudy (mertsj):

Did you find c yet?

OpenStudy (magbak):

yes it is 32

OpenStudy (mertsj):

That is c^2

OpenStudy (magbak):

c=5.656854249

OpenStudy (magbak):

@Mertsj

OpenStudy (magbak):

The answer is D

OpenStudy (mertsj):

Divide that by a

OpenStudy (mertsj):

yep

OpenStudy (magbak):

Ok I have 2 more question can you PLEASE CAN YOU CONTINUE TO HELP ME

OpenStudy (mertsj):

I can't. Have to go to bed. Have to get up at 5 am. Sorry

OpenStudy (magbak):

wa IT IS A VERY QUICK QUESTION